OpenStudy (schrodinger):

I actually understand how to do this, but the problem was so atrocious that my teacher threw it out and said, "Don't worry about it. Just forget it." Anyone want to take a dig at it simply for the sake of doing so? :P Find the appropriate value for c to make vectors v and w parallel. v = {5, 4} w = {4, c}

OpenStudy (anonymous):

ou are using 0.05 gallons per mile of a SUV gas tank that holds 16 gallons. Make a T-table for the equation for this problem, using the x-values 0, 160, 320. Type in the ordered pairs you get as your answer

OpenStudy (phi):



OpenStudy (anonymous):

so what is the anwser

OpenStudy (schrodinger):

@phi is on the right track, but not there. :3

OpenStudy (schrodinger):

As in, \[A*B = ||A||*||B||\cos \theta\]needs to be applied.

OpenStudy (anonymous):

w = {4, 3.2}

OpenStudy (schrodinger):

Yup.